# A non-uniform metal distribution in the ram-pressure stripped M86 group of galaxies

### v1

@@ Abstract
The chemical enrichment of X-ray-emitting hot halos has primarily been studied in closed-box galaxy clusters. Investigating the metal content of lower-mass, open systems can serve as a valuable tracer for understanding their dynamical history and the extent of chemical enrichment mechanisms in the Universe. In this context, we use an 85.6 ks XMM-Newton observation to study the spatial distribution of the abundance ratios of Mg, Si, and S with respect to Fe in the hot gas of the ram-pressure-stripped M86, which has undergone morphological transformations. We report that the chemical composition in the M86 galaxy core is more similar to the rest of the hot gaseous content of the Universe than to its stellar population. This result indicate that even extreme supersonic ram-pressure is insufficient to strip the inner part of a galaxy of its hot atmosphere. Comparison with other galaxies undergoing ram-pressure stripping suggests that stripping the "primordial" atmosphere of a galaxy requires a combination of ram-pressure stripping and strong radio-mechanical AGN activity. The X-ray emission structures within M86, the plume and the tail, are found to be relatively isothermal. We observe that the Mg/Fe ratio in the plume is higher than in the M86 galaxy core and is consistent with that in the M86 group outskirts and the Virgo ICM, suggesting that the plume might originate from the low-entropy group gas due to a galaxy-galaxy collision rather than from the ram-pressure stripping of the dense galaxy core.
@@
M86 is a blue-shifted giant elliptical galaxy falling into the Virgo Cluster from the far side with a relative velocity 1500 km with respect to Virgo Cluster central galaxy M87 (Finoguenov et al., 2004). Furthermore, the highly extended hot X-ray emitting halo of M86 has a shallower X-ray surface brightness profile and larger gas mass fraction, which resembles a group of galaxies (see, e.g., Böhringer et al. (1994) and references therein). Additionally, Binggeli et al. (1993) has shown the blue-shifted tail of M86 consists dwarf ellipticals, which implies M86 galaxy is in fact at the core of a larger structure, the M86 group of galaxies. Within the M86 group, there is an X-ray emitting plume at 3 arcmin north of the M86 galaxy and extending more than projected 10 arcmin to northwest, first reported by Forman et al. (1979) with Einstein observations. Furthermore, using ROSAT PSPC and HRI data, Rangarajan et al. (1995) identified another X-ray structure, a north-eastern arm extending from the M86 galaxy core to at least 5 arcmin to north-east. Chandra analysis by Randall et al. (2008) further revealed a long tail of emission, extending from the plume to northwest with a true length of 380 kpc. In that study they also identified two temperature structures, a cooler component of 0.8 keV that can be associated with the M86 galaxy halo, and a hotter component with 1.2 keV associated with the M86 group halo.
@@ 2 Observation and data analysis
M86 was observed with a single XMM-Newton observation on 1 July 2002, with ObsID 0108260201. The total exposure time is 85.6 ks and the net exposure time for each instrument is presented in Table 1. For data reduction, we used the XMM-Newton Science Analysis System (SAS) version 21.0.0, along with the integrated Extended Source Analysis Software (ESAS v9.0111https://www.cosmos.esa.int/web/xmm-newton/xmm-esas) package. For point source detection, we used Chandra CIAO 4.15.
@@ 2.1 EPIC
The calibrated photon event files for MOS and pn data were created using emchain and epchain, respectively. Further, we filtered the event files to exclude soft-proton (SP) flares by building Good Time Interval (GTI) files using the routine espfilt, where we created light curves for the full-band spectrum in 100s bins. After calculating the best-fitting mean count rate for the Gaussian distribution, , and the standard deviation, , we rejected time bins outside the interval . Point source detection and exclusion were performed using the CIAO algorithm wavdetect. We processed the XMM–Newton/EPIC data using mosspectra and pnspectra, which call the evselect, rmfgen, and arfgen tasks to extract spectra and generate RMFs and ARFs, respectively. We kept single, double, triple, and quadruple events in the MOS data (pattern12) and single events (pattern=0) for pn.
@@ 2.1.1 Background modelling approach
We model the hot gas of M86 with a two-temperature thermal emission model (vapec + vapec), where the abundance ratios of N, O, and Ne with respect to Fe are fixed to the values obtained from the RGS results. The temperatures and the abundance values of Mg, Si, S, and Fe are left free. In this EPIC FOV + RASS fit for determining the background parameters, we find that the dominant cooler central gas of M86 is at keV, while the extended halo is at keV, both in line with previous measurements (Randall et al., 2008; Mernier et al., 2016a). The methods for estimating all background components and the use of the RASS data are discussed in detail in the following subsections.
@@ 2.1.5 Local Hot Bubble, Galactic Thermal Emission and Unresolved Point Sources
The Milky Way originated X-ray foregrounds include two primary components. The first is the Local Hot Bubble (LHB), a local (i.e., unabsorbed) shock-heated region between the heliosphere and the local interstellar medium (Kuntz & Snowden, 2008). The LHB is modeled using an unabsorbed single-temperature thermal emission model (apec) with Solar abundances, and its expected temperature is approximately . The second Galactic foreground component is the Galactic Thermal Emission (GTE), which originates from the X-ray thermal emission of the Milky Way halo. The GTE lies beyond most of the absorption caused by the Galactic ISM and is therefore represented using an absorbed thermal emission model (phabs*apec) with Solar abundances. The hydrogen column density for the GTE absorption is set to (Mernier et al., 2015). The temperature of the GTE component is expected to vary between across the sky (Das et al., 2019). Although the mean GTE temperature is approximately (McCammon et al., 2002), recent studies have reported higher temperatures in the sky vicinity of M86 (e.g., 0.48 keV for the GTE in the Virgo Cluster; Gatuzz et al., 2023). Therefore, we expect a higher GTE temperature. Another background component is the Unresolved Point Sources (UPS), which corresponds to the integrated X-ray emission from active galactic nuclei (AGNs), dominant in the hard band, and galaxies, which contribute most of the UPS counts in the soft band (Lehmer et al., 2012). We model the UPS emission using a power-law model with a photon index of (De Luca & Molendi, 2004).
To estimate the background components mentioned above, we fit the FOV EPIC spectra combined with RASS spectra, which is more sensitive to the soft band. Because the M86 X-ray emission is highly extended and there are X-ray bright galaxies in the vicinity of M86, we were unable to extract a RASS spectra from an annulus surrounding M86. Therefore we moved our extraction pointing 1° to North and 1° to West from the M86 center and extracted a circular region with 1° radius in Galactic coordinates. Based on previous studies reporting a significant drop in the Virgo ICM surface brightness profile at distances similar to our RASS extraction region (Simionescu et al., 2017), we assume the Virgo ICM contribution is negligible in the RASS spectra. To confirm that our RASS extraction region is not affected by any contamination, we carefully checked the diffuse X-ray maps from HEASARC X-Ray Background Tool222https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/xraybg/xraybg.pl. In the fitting, we coupled all the LHB, GTE and UPS model parameters in EPIC with their counterpart in RASS, with a free scaling parameter. We found LHB temperature to be keV, and the GTE temperature to be keV, presented in Table 5.
@@ 2.1.7 EPIC region selection and spectral modelling
In fitting each extracted spectrum, we model all background components with fixed parameters obtained from the FOV fit, including their normalizations. Afterward, we scaled down the background normalizations for each region and corrected for vignetting effects using exposure maps for each instrument (MOS1, MOS2, pn) and region.
For the region number 4 and 3, which corresponds to the galaxy core and the surrounding circular region; and in the North-Eastern Arm, which is region number 5, we observed an additional background component corresponding to the integrated X-ray emission from Low-Mass X-ray Binaries (LMXBs) in M86. To model this component, we used a redshifted power-law model with a fixed photon index (Irwin et al., 2003; Su et al., 2017). Initially, we fixed the normalization ratio of this component to that of the core hot gas to the value we observed in Kara et al. (2024). Later, we allowed its normalization to vary. We note that LMXB component could not be constrained in the FOV fit and other regions. In the regions with the prominent LMXB emission, we also allowed the normalization of the other astrophysical power-law model component accounting for the UPS to vary. However, we observe that the change is not significant and causes very little changes in abundance ratios (less than 5%), so we fixed the parameter to the original scaled value. We note that the background parameters except LMXB were fixed for all fits across all regions, and the normalizations for each region were scaled down separately.
@@ 2.2 RGS
We used the SAS task rgsproc to process the RGS data. Since the RGS and MOS cameras are co-aligned on the satellite, the RGS 1 and RGS 2 data were filtered for flaring events using the same GTIs applied to MOS 1 and MOS 2, respectively. The RGS spectra were extracted using 90% of the pulse height distribution.
@@ 2.2.1 Selecting the RGS region size
First, we extracted the spectrum from (i) 90% of the point spread function (PSF) using rgsproc (xpsfincl=90), which roughly corresponds to a 0.8 arcmin region. Additionally, we extracted a broader region from (ii) 99% of the PSF (xpsfincl=99), corresponding to a 3.4 arcmin region. Due to its larger size, the second spectrum is expected to exhibit more non-intrinsic line broadening and to include emissions from a larger area.
@@ 2.2.2 Selecting the RGS background spectra
We first created the (i) model background spectrum using the standard rgsbkgmodel routine. This background template file is based on the count rate in CCD 9 of the RGS, which is expected to contain no source counts. Additionally, we extracted (ii) a background spectrum from outside 98% of the PSF. This second background spectrum includes source counts from outside the galaxy core.
@@ 2.2.3 RGS spectral modelling
where is the spectral order and is the half energy width of the source in arcmin444See the XMM-Newton Users Handbook.. The broadening was corrected with the XSPEC convolution model rgsxsrc at the spectral fitting stage. We generated MOS1 image in the 0.46-1.55 keV (8–28 Å) band and the suffice brightness profile in the dispersion direction. We use first and second-order RGS spectra in 8–28 Å without combining R1 and R2 spectra. Similar to our EPIC analysis, we focus on vgadem model to estimate the multi-temperature structure of hot gas. The RGS spectral fit is presented in Fig. 3, where we combine R1 and R2 first-order spectra using rgscombine for visualization purposes. Additionally, to test the "O/Fe bias" observed in Kara et al. (2024), which is the discrepancy of O/Fe ratio between 1T and multi-temperature models, we also fit the RGS spectra with 1T (vapec) and 2T (vapec + vapec) models.
@@ 3.1 Galaxy core
We measure the elemental abundances in the hot halo pervading the galaxy core of M86 using both RGS and EPIC data. To improve consistency between the RGS and EPIC measurements, we extracted our core spectra from a rectangular region overlapping the RGS extraction band with a 0.8 arcsec width (Region 4), following a similar approach to Mernier et al. (2022). The elemental abundances in a CIE plasma are derived from the equivalent width (EW), which is proportional to the ratio of line flux to continuum flux. Due to imperfections in the effective areas of EPIC instruments, the continuum measurement may be biased (Mernier et al., 2015; Simionescu et al., 2019). Consequently, line fluxes may compensate for under- or overestimation of the continuum flux, leading to corresponding over- or underestimation of elemental abundances. We first performed a global fit using the vgadem model on the full-band EPIC spectra (0.55 keV–10 keV). Then, we repeated the global fitting separately for MOS (MOS1 + MOS2) and pn spectra. Additionally, we performed local fits of MOS and pn for narrow-band spectra corresponding to well-resolved elemental lines: 1–1.7 keV for Mg, 1.6–2.3 keV for Si, and 2.2–2.8 keV for S. In these more conservative local fits of MOS and pn, all parameters were frozen to the values derived from the global fit for that specific instrument, except for the normalization of the CIE model (vgadem) and the relevant elemental abundances, which were left free.
@@ 3.2 Radial distribution of metals
We measured abundances and abundance ratios of 8 regions, covering a projected region roughly 15 arcmin (80 kpc). For all regions, we used a multi-temperature vgadem model. Additionally, throughout this analysis we assume the Virgo ICM X-ray background does not change across the FOV. To test this assumption, we assumed the most extreme case where the ICM flux is set to zero and refit the outermost region, Tail 2 (Region number 8). We fixed the normalization of the ICM apec model to zero and let the other astrophysical scaled-down background normalizations to vary as a whole (with a multiplicative constant model). We see that for the case without the ICM model, the astrophysical background normalizations increase by 20%, while the abundance ratios change roughly by 5%. Since this fit provides poorer C-statistics, we decide to model the ICM emission for all regions with a scaled-down fixed normalizations according to the vignetting-corrected size of that region, as described in Sect 2.1.1.
@@
In the following sections, we compare our results for each abundance ratio with the average ratio in hot gas obtained from the CHEERS (CHEmical Evolution RGS Sample) sample (Mernier et al., 2018c), the ratios in the stars of elliptical galaxies, and the Virgo ICM ratios from Simionescu et al. (2015), which represent the chemical composition of the ambient medium through which M86 travels.

### v2

# A non-uniform metal distribution in the ram-pressure stripped circumgalactic medium of M86

@@ Abstract
The chemical enrichment of X-ray-emitting hot halos has primarily been studied in closed-box galaxy clusters. Investigating the metal content of lower-mass, open systems can serve as a valuable tracer for understanding their dynamical history and the extent of chemical enrichment mechanisms in the Universe. In this context, we use an 85.6 ks XMM-Newton observation to study the spatial distribution of the abundance ratios of Mg, Si, and S with respect to Fe in the hot gas of the ram-pressure-stripped M86, which has undergone morphological transformations. We report that the chemical composition in the M86 galaxy core is more similar to the rest of the hot gaseous content of the Universe than to its stellar population. This result indicates that even supersonic ram-pressure is insufficient to strip the inner part of a galaxy of its hot atmosphere. Comparison with other galaxies undergoing ram-pressure stripping suggests that stripping the "primordial" atmosphere of a galaxy requires a combination of ram-pressure stripping and strong radio-mechanical AGN activity. The X-ray emission structures within M86, the plume and the tail, are found to be relatively isothermal. We observe that the Mg/Fe ratio in the plume is higher than in the M86 galaxy core and is consistent with that in the M86 group outskirts and the Virgo ICM, suggesting that the plume might originate from the low-entropy outer gas rather than from the recent ram-pressure stripping of the dense galaxy core.
@@
M86 is a blue-shifted giant elliptical galaxy falling into the Virgo Cluster from the far side with a relative velocity 1500 km with respect to Virgo Cluster central galaxy M87 (Finoguenov et al., 2004). Furthermore, the highly extended hot X-ray emitting halo of M86 has a shallower X-ray surface brightness profile and larger gas mass fraction, which resembles a group of galaxies (see, e.g., Böhringer et al. (1994) and references therein). Additionally, Binggeli et al. (1993) has shown the blue-shifted tail of M86 consists dwarf ellipticals, which implies M86 galaxy is in fact at the core of a larger structure, the M86 group of galaxies. Within the M86 group, there is an X-ray emitting plume at 3 arcmin north of the M86 galaxy and extending more than projected 10 arcmin to northwest, first reported by Forman et al. (1979) with Einstein observations. Furthermore, using ROSAT PSPC and HRI data, Rangarajan et al. (1995) identified another X-ray structure, a north-eastern arm extending from the M86 galaxy core to at least 5 arcmin to north-east. Chandra analysis by Randall et al. (2008) further revealed a long tail of emission, extending from the plume to northwest with a true length of 380 kpc. In that study they also identified two temperature structures, a cooler component of 0.8 keV that can be associated with the M86 galaxy halo, and a hotter component with 1.2 keV associated with the M86 group halo. The temperature maps of M86 show smooth gradients with no discontinuities at the apparent boundaries (Finoguenov et al., 2004; Randall et al., 2008). Therefore, these structures are not foreground or background objects, but are truly bound to the same system.
@@ 2 Observation and data analysis
M86 was observed with a single XMM-Newton observation on 1 July 2002, with ObsID 0108260201. The total exposure time is 85.6 ks and the net exposure time for each instrument is presented in Table 1. For data reduction, we use the XMM-Newton Science Analysis System (SAS) version 21.0.0, along with the integrated Extended Source Analysis Software (ESAS v9.0111https://www.cosmos.esa.int/web/xmm-newton/xmm-esas) package. For point source detection, we use Chandra CIAO 4.15.
@@ 2.1 EPIC
The calibrated photon event files for MOS and pn data are created using emchain and epchain, respectively. Further, we filter the event files to exclude soft-proton (SP) flares by building Good Time Interval (GTI) files using the routine espfilt, where we create light curves for the full-band spectrum in 100s bins. After calculating the best-fitting mean count rate for the Gaussian distribution, , and the standard deviation, , we reject time bins outside the interval . Point source detection and exclusion are performed using the CIAO algorithm wavdetect. We process the XMM–Newton/EPIC data using mosspectra and pnspectra, which call the evselect, rmfgen, and arfgen tasks to extract spectra and generate RMFs and ARFs, respectively. We keep single, double, triple, and quadruple events in the MOS data (pattern12) and single events (pattern=0) for pn.
@@ 2.1.1 Background modelling approach
We model the hot gas of M86 with a two-temperature thermal emission model (vapec + vapec), where the abundance ratios of N, O, and Ne with respect to Fe are fixed to the values obtained from the RGS results (See Sect. 3.1). The temperatures and the abundance values of Mg, Si, S, and Fe are left free. In this EPIC FOV + RASS fit for determining the background parameters, we find that the dominant cooler central gas of M86 is at keV, while the extended halo is at keV, both in line with previous measurements (Randall et al., 2008; Mernier et al., 2016a). The methods for estimating all background components and the use of the RASS data are discussed in detail in the following subsections.
@@ 2.1.5 Local Hot Bubble, Galactic Thermal Emission and Unresolved Point Sources
The Milky Way originated X-ray foregrounds include two primary components. The first is the Local Hot Bubble (LHB), a local (i.e., unabsorbed) shock-heated region between the heliosphere and the local interstellar medium (Kuntz & Snowden, 2008). The LHB is modeled using an unabsorbed single-temperature thermal emission model (apec) with Solar abundances, and its expected temperature is approximately . The second Galactic foreground component is the Galactic Thermal Emission (GTE), which originates from the X-ray thermal emission of the Milky Way halo. The GTE lies beyond most of the absorption caused by the Galactic ISM and is therefore represented using an absorbed thermal emission model (phabs*apec) with Solar abundances. The hydrogen column density for the GTE absorption is set to (Mernier et al., 2015). The temperature of the GTE component is expected to vary between across the sky (Das et al., 2019). Although the mean GTE temperature is approximately (McCammon et al., 2002), recent studies have reported higher temperatures in the sky vicinity of M86 (e.g., 0.48 keV for the GTE in the Virgo Cluster; Gatuzz et al., 2023b). Therefore, we expect a higher GTE temperature. Another background component is the Unresolved Point Sources (UPS), which corresponds to the integrated X-ray emission from active galactic nuclei (AGNs), dominant in the hard band, and galaxies, which contribute most of the UPS counts in the soft band (Lehmer et al., 2012). We model the UPS emission using a power-law model with a photon index of (De Luca & Molendi, 2004).
To estimate the background components mentioned above, we fit the FOV EPIC spectra combined with RASS spectra, which is more sensitive to the soft band. Because the M86 X-ray emission is highly extended and there are X-ray bright galaxies in the vicinity of M86, we were unable to extract a RASS spectra from an annulus surrounding M86. Therefore we moved our extraction pointing 1° to North and 1° to West from the M86 center and extracted a circular region with 1° radius in Galactic coordinates. Based on previous studies reporting a significant drop in the Virgo ICM surface brightness profile at distances similar to our RASS extraction region (Simionescu et al., 2017), we assume the Virgo ICM contribution is negligible in the RASS spectra. To confirm that our RASS extraction region is not affected by any contamination, we carefully checked the diffuse X-ray maps from HEASARC X-Ray Background Tool222https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/xraybg/xraybg.pl. In the fitting, we couple all the LHB, GTE and UPS model parameters in EPIC with their counterpart in RASS, with a free scaling parameter. We find LHB temperature to be keV, and the GTE temperature to be keV, presented in Table 5.
@@ 2.1.7 EPIC region selection and spectral modelling
In fitting each extracted spectrum, we model all background components with fixed parameters obtained from the FOV fit, including their normalizations. Afterward, we scale down the background normalizations for each region and corrected for vignetting effects using exposure maps for each instrument (MOS1, MOS2, pn) and region.
For the region number 4 and 3, which corresponds to the galaxy core and the surrounding circular region; and in the North-Eastern Arm, which is region number 5, we observe an additional background component corresponding to the integrated X-ray emission from Low-Mass X-ray Binaries (LMXBs) in M86. To model this component, we use a redshifted power-law model with a fixed photon index (Irwin et al., 2003; Su et al., 2017). Initially, we fix the normalization ratio of this component to that of the core hot gas to the value we observed in Kara et al. (2024). Later, we allow its normalization to vary. We note that LMXB component cannot be constrained in the FOV fit and other regions. In the regions with the prominent LMXB emission, we also allow the normalization of the other astrophysical power-law model component accounting for the UPS to vary. However, we observe that the change is not significant and causes very little changes in abundance ratios (less than 5%), so we fix the parameter to the original scaled value. We note that the background parameters except LMXB are fixed for all fits across all regions, and the normalizations for each region are scaled down separately.
@@ 2.2 RGS
We use the SAS task rgsproc to process the RGS data. Since the RGS and MOS cameras are co-aligned on the satellite, the RGS 1 and RGS 2 data are filtered for flaring events using the same GTIs applied to MOS 1 and MOS 2, respectively. The RGS spectra are extracted using 90% of the pulse height distribution.
@@ 2.2.1 Selecting the RGS region size
First, we extract the spectrum from (i) 90% of the point spread function (PSF) using rgsproc (xpsfincl=90), which roughly corresponds to a 0.8 arcmin region. Additionally, we extract a broader region from (ii) 99% of the PSF (xpsfincl=99), corresponding to a 3.4 arcmin region. Due to its larger size, the second spectrum is expected to exhibit more non-intrinsic line broadening and to include emissions from a larger area.
@@ 2.2.2 Selecting the RGS background spectra
We first create the (i) model background spectrum using the standard rgsbkgmodel routine. This background template file is based on the count rate in CCD 9 of the RGS, which is expected to contain no source counts. Additionally, we extract (ii) a background spectrum from outside 98% of the PSF. This second background spectrum includes source counts from outside the galaxy core.
@@ 2.2.3 RGS spectral modelling
where is the spectral order and is the half energy width of the source in arcmin444See the XMM-Newton Users Handbook.. The broadening is corrected with the XSPEC convolution model rgsxsrc at the spectral fitting stage. We generate MOS1 image in the 0.46-1.55 keV (8–28 Å) band and the suffice brightness profile in the dispersion direction. We use first and second-order RGS spectra in 8–28 Å without combining R1 and R2 spectra. Similar to our EPIC analysis, we focus on vgadem model to estimate the multi-temperature structure of hot gas. The RGS spectral fit is presented in Fig. 3, where we combine R1 and R2 first-order spectra using rgscombine for visualization purposes. Additionally, to test the "O/Fe bias" observed in Kara et al. (2024), which is the discrepancy of O/Fe ratio between 1T and multi-temperature models, we also fit the RGS spectra with 1T (vapec) and 2T (vapec + vapec) models.
@@ 3.1 Galaxy core
We measure the elemental abundances in the hot halo pervading the galaxy core of M86 using both RGS and EPIC data. To improve consistency between the RGS and EPIC measurements, we extract our core spectra from a rectangular region overlapping the RGS extraction band with a 0.8 arcsec width (Region 4), following a similar approach to Mernier et al. (2022). The elemental abundances in a CIE plasma are derived from the equivalent width (EW), which is proportional to the ratio of line flux to continuum flux. Due to imperfections in the effective areas of EPIC instruments, the continuum measurement may be biased (Mernier et al., 2015; Simionescu et al., 2019). Consequently, line fluxes may compensate for under- or overestimation of the continuum flux, leading to corresponding over- or underestimation of elemental abundances. We first perform a global fit using the vgadem model on the full-band EPIC spectra (0.55 keV–10 keV). Then, we repeat the global fitting separately for MOS (MOS1 + MOS2) and pn spectra. Additionally, we perform local fits of MOS and pn for narrow-band spectra corresponding to well-resolved elemental lines: 1–1.7 keV for Mg, 1.6–2.3 keV for Si, and 2.2–2.8 keV for S. In these more conservative local fits of MOS and pn, all parameters are frozen to the values derived from the global fit for that specific instrument, except for the normalization of the CIE model (vgadem) and the relevant elemental abundances, which are left free.
@@ 3.2 Radial distribution of metals
We measure abundances and abundance ratios of 8 regions, covering a projected region roughly 15 arcmin (80 kpc). For all regions, we use a multi-temperature vgadem model. Additionally, throughout this analysis we consider the Virgo ICM X-ray background to be constant across the FOV. To test the validity of this assumption, we evaluate the most extreme scenario by setting the ICM flux to zero and refitting the outermost region, Tail 2 (Region number 8). To perform this test, we fix the normalization of the ICM apec model to zero and let the other astrophysical scaled-down background normalizations to vary as a whole (with a multiplicative constant model). We see that for the case without the ICM model, the astrophysical background normalizations increase by 20%, while the abundance ratios change roughly by 5%. Since this fit provides poorer C-statistics, we decide to model the ICM emission for all regions with a scaled-down fixed normalizations according to the vignetting-corrected size of that region, as described in Sect 2.1.1.
@@
In the following sections, we compare our results for each abundance ratio with the average ratio in hot gas obtained from the CHEERS (CHEmical Evolution RGS Sample) sample (Mernier et al., 2018b), the ratios in the stars of elliptical galaxies (Conroy et al., 2013), and the Virgo ICM ratios from Simionescu et al. (2015), which represent the chemical composition of the ambient medium through which M86 travels.